# High-throughput Parasitic-independent Probe Thermal Resistance Calibration for Robust Thermal Mapping with Scanning Thermal Microscopy

### v3

@@ VI Conclusion
We experimentally quantified the thermal resistance of the SThM probe using a circuit-based calibration method and employed FEM modeling to investigate the underlying heat transfer mechanisms. The contributions of various probe–sample heat transfer channels, along with their corresponding measurement strategies, were systematically examined. Furthermore, the thermal conductivity of a 15 nm Al thin film, deposited by e-beam evaporation on a few-micrometer-thick SiO2 substrate, is measured. The robustness of the pixel-to-pixel correlation approach for extracting representative thermal signals is discussed. In addition, statistical and measurement uncertainties in thermal resistance are quantified, considering the underlying assumptions. The measured thermal conductivity of the Al thin film is in good agreement with the theoretical framework, with the slight variation attributed to scattering at the Al/SiO2 interface.

### v4

@@ V Conclusion
We experimentally quantified the thermal resistance of the SThM probe using a circuit-based calibration method and employed FEM modeling to investigate the underlying heat transfer mechanisms. The contributions of various probe–sample heat transfer channels, along with their corresponding measurement strategies, were systematically examined. Furthermore, the thermal conductivity of a 15 nm Al thin film, deposited by e-beam evaporation on a few-micrometer-thick SiO2 substrate, is measured. The robustness of the pixel-to-pixel correlation approach for extracting representative thermal signals is discussed. In addition, statistical and measurement uncertainties in thermal resistance are quantified, considering the underlying assumptions. The measured thermal conductivity of the Al thin film is in good agreement with the theoretical frameworks and the published experimental data.